OpenStudy (anonymous):

Which of the following equations has solutions of 3 ± square root of 5 ? (x - 5)2 = 3 (x - 3)2 = 5 (x + 3)2 = 5 (x + 5)2 = 3

OpenStudy (callisto):

For (x - 5)^2 = 3 x-5 = ± sqrt3 x = 5 ± sqrt 3 For (x - 3)^2 = 5 x -3 = ± sqrt 5 x = 3 ± sqrt5 For (x + 3)^2 = 5 (x + 3) = ± sqrt5 x = -3 ± sqrt5 For (x + 5)^2 = 3 x + 5 = ± sqrt3 x = -5 ± sqrt 3 So, what's the answer?

OpenStudy (anonymous):

second one

OpenStudy (callisto):

Yes. Hope you know how to do it next time :|